A boy cycles x km at y km per hour and p km at q km per hour.Find his average speed for the whole journey.
Answer by mananth(16949)

A boy cycles x km at y km per hour and p km at q km per hour.Find his average speed for the whole journey.
..
time taken = x/y
also
time taken = p/q
..
Total time = x/y + p/q
= xq+py /qy
Total distance = x+p km
Average speed = total distance / total time
x+p / xq+py/qy
(x+p)*qy/ (xq+py)km/hour